OpenStudy (anonymous):

A four digit number (with no repetitions) is to be formed from the set of numbers {0,1,2,3,4,5,6,7}. Find the probability that the number a) is less than 4000 b) is less than 4000, given that it is greater than 3000

OpenStudy (nali):

i have the answer to part b only numbers greater than 3000 can start with 3, 4, 5, 6 or 7 numbers from the above sample space less than 4000 have to start with 3 reqd. probability = 1/5 = .2 and here is my source ' http://au.answers.yahoo.com/question/index?qid=20100403031552AAYdtuw

OpenStudy (nali):

for the first part @JayDS

OpenStudy (nali):

there are 8 numbers form 0-7 but we have only four places without repetetions

OpenStudy (nali):

so if for the first place there are 8 choices for the second place there are 7 choices for the third place there are 6 choices for the forth place there are 5 choices

OpenStudy (nali):

so to get the total options u multiply 8x7x6x5

OpenStudy (anonymous):

yep kk.

OpenStudy (anonymous):

but probability so it would be over the total

OpenStudy (anonymous):

so 8x7x6x5/8! right?

OpenStudy (nali):

then to find the probability of numbers less than 4000 u look at the numbers that can be formed that are less than 4000 which will be mostly 3765 and under

OpenStudy (nali):

meaning that we are going to be using numbers 3,2,1,0 for the first place and there will be 7 numbers left for the second place and 6 for the third and 5 for the forth

OpenStudy (nali):

so it will be 4x7x6x5/ 8x7x6x5

OpenStudy (nali):

@JayDS by the way the answer that is for part b is wrong

OpenStudy (nali):

it is supposed to be the follwoing

OpenStudy (nali):

if the numbers have to be bigger than 3000 and less than 4000 then for the first place there is only one number right which is 3

OpenStudy (nali):

for the second is 7 for the third is 6 and for the forth is 5

OpenStudy (nali):

so the answer for part b would be 1x7x6x5 / 8x7x6x5

OpenStudy (nali):

@JayDS

OpenStudy (anonymous):

but the answer in my textbook says 0.2

OpenStudy (anonymous):

for part b.

OpenStudy (nali):

ohh ok then what does it say for part a

OpenStudy (anonymous):

3/7=0.429

OpenStudy (nali):

oh ok then sorry all my answers are wrong :(